# DiP: A Scalable, Energy-Efficient Systolic Array for Matrix Multiplication Acceleration

### v1

@@ IV-C DiP Evaluation
DiP is assessed against a TPU-like (WS-based) architecture using transformer workloads. Nine widely used transformer models are chosen to span various application domains and represent a spectrum of model sizes, from small language models (SLMs) to large language models (LLMs). These models organized in three types: Encoder-Decoder, Encoder-only, and Decoder-only. Encoder-Decoder models include Vanilla Transformer [10], T5 [30], and BART [31]; Encoder-only models include BERT [35], ALBERT [36], and Transformer-XL [37]; while Decoder-only models include GPT-2 [32], GPT-3 [33], and LLaMA [34]. The models are selected with hyper-parameters to cover a diverse range of workloads. Sequence lengths are chosen from a range of 64 to 2048, as (64, 128, 256, 512, 1024, 2048). Additionally, model’s hidden size ($d_{\text{model}}$) varies across (512, 768, 1024, 1280, 5120), while head size ($d_{k}$) is set to either (64, 128). Moreover, FFN size ($d_{\text{FFN}}$) is configured with values of (2048, 3072, 4096, 5120).
Figure 6: Evaluation of DiP and TPU-like architecture at size 64x64 with MHA and FFN transformer’s workloads. The evaluation includes actual energy consumption (a, b) and latency (c, d) across various workloads dimensions of matrix multiplication (M-N-K).
TABLE IV: Comparison with Other Accelerators
DiP Google TPU[23] Groq ThinkFast TSP[38] Alibaba Hanguang 800[39]
Architecture 64$\times$64,
4,096 MACs
256$\times$256, 65,536 MACs Tensor Stream Processor Tensor Cores
Max Frequency 1GHz 700MHz 900MHz 700MHz
Precision INT8 INT8 INT8, FP16 INT8, INT16, FP24
Technology 22nm 28nm 14nm 12nm
Power (W) 0.858 40–50 300 275.9
Area (mm²) 1 200 725 709
Peak Performance (TOPS) 8.2 92 820 825
Norm. Perform. (TOPS)1 8.2 5.75
Area Norm. Perform. (TOPS/mm²)2,3 8.2 0.46 0.411 0.423
Energy Efficiency (TOPS/W)3 9.55 2.15 2.73 2.99
1Normalized Peak Performance at systolic array size of 64$\times$64
2Normalized Peak Performance to die area (mm²)
3Power and Area are normalized to 22nm using DeepScaleTool [40]
@@
DiP and TPU-like architectures, each with a size of 64x64, are used for the evaluation. This architecture size aligns well with matrix tiling, as the head size for most transformer models is either 64 or 128. Cycle-accurate simulations are performed to evaluate both DiP and TPU-like implementations in terms of actual latency for each workload and energy consumption. Matrix tiling is used to process matrix multiplication workloads on DiP and TPU-like architectures by dividing the input matrices $M_{1}$ and $M_{2}$ into sub-matrices (tiles) of 64x64. By studying many transformer models, the majority of MHA and FFN workload dimensions are divisible by 64. The multiplication is performed per tile as follows; For DiP and TPU-like architectures, every tile of $M_{2}$ is loaded once and remains stationary throughout the computation for the corresponding output tile. For each tile of $M_{2}$, respective tiles from $M_{1}$ is iteratively loaded, multiplied, and saved as output partial summation (psum) tiles. After processing all tiles, the final output matrix $O$ is constructed by accumulating the associated psum tiles.
@@
Figure 6 compares the energy consumption and latency of DiP and TPU-like 64x64 architectures for MHA and FFN workloads across varying dimensions (M-N-K). DiP consistently outperforms TPU-like implementation for MHA and FFN workloads. Energy improvements for MHA workloads range from 1.81x for smaller workloads to 1.25x for larger ones, while FFN workloads show a similar trend with improvements from 1.8x to 1.25x. These results highlight the energy-efficiency of DiP. Additionally, actual latency for MHA and FFN workloads demonstrates DiP’s performance against TPU-like implementation, offering up to 1.49x improvements for smaller workloads, gradually reducing to approximately 1.03x for larger workloads. The breakdown of latency improvement happens for larger workloads as TPU-like architecture hides the latency associated with loading more $M_{1}$ tiles per every new $M_{2}$ tile. In contrast, for small to medium-sized workloads, TPU-like architectures incur the TFPU penalty of loading $M_{1}$ tile associated with each new $M_{2}$ tile. Additionally, TPU-like architectures still face the overhead of Input/Output FIFOs, which impact power consumption, latency, and TFPU. The evaluation highlights DiP as an energy-efficient design, making it a compelling alternative to TPU-like architectures.
IV-D Comparison with Related Work
@@
Table IV compares DiP architecture with Google TPU [23], Groq ThinkFast TSP [38], and Alibaba Hanguang 800 [39], highlighting DiP’s performance and energy efficiency. The DiP architecture features 4,096 MACs in a 64×64 configuration, operating at 1 GHz with INT8 precision on 22nm technology. As each accelerator is implemented in different technology, the performance metrics are normalized to 22nm using DeepScaleTool [40]. The proposed DiP achieves significant energy efficiency, reaching 9.55 TOPS/W, and delivers a normalized performance per area of 8.2 TOPS. This demonstrates its capability to provide high computational throughput while minimizing energy consumption, showcasing its compact and optimized design. These metrics make DiP particularly well-suited for energy-efficient Transformers-based applications.
@@ V Conclusion
In this paper, a diagonal-input and permutated weight-stationary (DiP) systolic array is proposed to accelerate matrix multiplication. DiP features an architecture of NxN PEs, where each PE performs Multiply-Accumulate operations. DiP adopts a novel dataflow that eliminates the input and output synchronization FIFO buffers required by the conventional WS systolic array. As a result, DiP outperforms WS across all metrics, including throughput, latency, area, and power consumption. Additionally, the analytical models for latency, throughput, FIFO overhead, and TFPU are developed for DiP and WS architectures. The proposed DiP architecture outperforms the WS counterparts in terms of throughput by up to 50%, and TFPU by up to 50%. Moreover, hardware design space exploration is presented for both DiP and WS architectures using commercial 22nm technology, demonstrating power consumption savings of up to 19.95%, area savings of up to 8.12% at 1 GHz, and energy efficiency per area improvement of up to 2.02x. Furthermore, DiP is evaluated using various transformer workloads from widely-used models such as GPT-2, GPT-3, BERT, BART, and LLaMA. DiP outperforms TPU-like architecture, achieving energy improvements ranging from 1.25x to 1.81x and latency improvements ranging from 1.03x to 1.49x across various transformer’s MHA and FFN workloads. A comparison between relevant accelerators and DiP is discussed, achieving performance of 8.2 TOPS and energy efficiency of 9.55 TOPS/W, which is promising for energy-efficient transformer-based applications. This paper serves as the foundation for DiP architecture and dataflow. Future extensions aim to scale the architecture and explore sparsity in transformers, which will further enhance energy efficiency and acceleration rates.

### v2

@@ IV-C DiP Evaluation
DiP is evaluated against a TPU-like (WS-based) architecture using transformer workloads. Nine widely used transformer models are selected to span various application domains and represent a spectrum of model sizes, from small language models (SLMs) to large language models (LLMs). These models organized in three types: Encoder-Decoder, Encoder-only, and Decoder-only. Encoder-Decoder models include Vanilla Transformer [10], T5 [30], and BART [31]; Encoder-only models include BERT [35], ALBERT [36], and Transformer-XL [37]; while Decoder-only models include GPT-2 [32], GPT-3 [33], and LLaMA [34]. The models are selected with hyper-parameters to cover a diverse range of workloads. Sequence lengths are chosen from a range of 64 to 2048, specifically (64, 128, 256, 512, 1024, 2048). Additionally, the model’s hidden size ($d_{\text{model}}$) varies across (512, 768, 1024, 1280, 5120), while the head size ($d_{k}$) is set to either (64, 128). Moreover, the FFN size ($d_{\text{FFN}}$) is configured with values of (2048, 3072, 4096, 5120).
TABLE III: Comparison with State-of-The-Art Accelerators
  DiP
(this work)
Google TPU V4i[25] Groq TSP
[38]
 OPTIMUS
[39]
DTATrans [40]   Eidetic
[41]
Architecture 64$\times$64, MACs 4$\times$128$\times$128, MACs Tensor Processors SA-RCSC VSSA In-Memory Computing
Design Level Core-level, System-level, System-level, System-level, Core-level, Core-level,
Post-PnR Post-Silicon Post-Silicon Post-Syn Post-Syn Modeling
Frequency (GHz) 1 1.05 0.9 0.2 1 1.5
Precision INT8 INT8 INT8 INT16 Multiple Multiple
Technology 22nm 7nm 14nm 28nm 40nm 22nm
Power (W) 0.858 175 300 0.732 0.803 235.7
Area (mm2) 1 400 725 5.2 1.49 266.4
Peak Throughput (TOPS) 8.192 138 820 0.5 1.304 175.5
Area Efficiency (TOPS/mm²) 8.192 0.345 1.131 0.096 0.979 0.659
Energy Efficiency (TOPS/W) 9.548 0.786 2.733 0.683 1.623 0.745
Area Efficiency (TOPS/mm2)(1) 8.192 0.017 0.412 0.153 2.984 0.659
Energy Efficiency (TOPS/W)(1) 9.548 0.345 1.823 0.797 2.470 0.745
(1)Power and area are normalized to 22nm using DeepScaleTool [42, 43]
@@
DiP and TPU-like architectures, each with a size of 64×64, are used for the evaluation. This architecture size aligns well with matrix tiling, as the head size for most transformer models is either 64 or 128. Cycle-accurate simulations are performed to evaluate both DiP and TPU-like implementations in terms of actual latency for each workload and energy consumption. Matrix tiling is used to process matrix multiplication workloads on DiP and TPU-like architectures by dividing the input matrices $M_{1}$ and $M_{2}$ into sub-matrices (tiles) of 64×64. By studying many transformer models, the majority of MHA and FFN workload dimensions are divisible by 64. The multiplication is performed per tile as follows; For DiP and TPU-like architectures, every tile of $M_{2}$ is loaded once and remains stationary throughout the computation for the corresponding output tile. For each tile of $M_{2}$, respective tiles from $M_{1}$ is iteratively loaded, multiplied, and saved as output psum tiles. After processing all tiles, the final output matrix is constructed by accumulating the associated psum tiles.
@@
Figure 5 compares the energy consumption and latency of DiP and TPU-like 64x64 architectures for MHA and FFN workloads across varying dimensions (M-N-K). DiP consistently outperforms TPU-like implementation for MHA and FFN workloads. Energy improvements for MHA workloads range from 1.81x for smaller workloads to 1.25x for larger ones, while FFN workloads show a similar trend with improvements from 1.8x to 1.25x. These results highlight the energy-efficiency of DiP. Additionally, actual latency for MHA and FFN workloads demonstrates DiP’s performance against TPU-like implementation, offering up to 1.49x improvements for smaller workloads, gradually reducing to approximately 1.03x for larger workloads. The breakdown of latency improvement happens for larger workloads as TPU-like architecture hides the latency associated with loading more $M_{1}$ tiles per every new $M_{2}$ tile. In contrast, for small to medium-sized workloads, TPU-like architectures incur the TFPU penalty of loading $M_{1}$ tile associated with each new $M_{2}$ tile. Additionally, TPU-like architectures still face the overhead of Input/Output synchronization FIFOs, which impact power consumption, latency, and TFPU. The evaluation highlights DiP as an energy-efficient design, making it a compelling alternative to TPU-like architectures.
IV-D Comparison with Related Accelerators
@@
Table III compares DiP with Google TPU V4i[25], Groq TSP[38], OPTIMUS [39], DTATrans[40], and Eidetic[41], highlighting DiP’s performance and energy efficiency for matrix multiplication acceleration. To ensure a fair and transparent comparison, Table III annotate whether each result corresponds to a system-level or a core-level implementation. Also, the maturity level for each design is annotated as follows: Google TPUv4i and Groq TSP represent post-silicon results; DiP is based on post-layout (post-PnR) implementation; OPTIMUS and DTATrans provide post-synthesis results; and Eidetic is based on modelling results. Since each accelerator is implemented in different technology, the performance metrics are normalized to 22nm using DeepScaleTool[42, 43]. Area and energy efficiencies results are presented before and after scaling to 22nm. In the case of Google TPUv4i, the Matrix Multiply Units (MXUs) occupy approximately 11% of the total chip area, yet deliver nearly the entire performance. While the power consumption of the MXUs alone is not explicitly reported, DiP provides higher area efficiency, achieving 8.192 TOPS/mm² compared to 3.14 TOPS/mm² and 0.157 TOPS/mm² for the Google TPU MXUs at 7nm and scaled 22nm, respectively. DiP architecture features 4,096 MACs in a 64×64 configuration, operating at 1 GHz with INT8 precision using commercial 22nm technology. At a 64×64 size with 4096 PEs, DiP offers peak throughput of 8.192 TOPS, energy efficiency with 9.548 TOPS/W, and delivers an area efficiency (computational density) with 8.192 TOPS/mm². These results demonstrate DiP’s capability to provide high throughput while minimizing energy consumption, showcasing its compact and optimized design. These metrics make DiP particularly well-suited for energy-efficient Transformers-based applications.
@@ V Conclusion
In this paper, a diagonal-input and permutated weight-stationary systolic array is proposed to accelerate matrix multiplication. DiP features an architecture of N×N PEs with a novel dataflow that eliminates the input and output synchronization FIFO buffers required by the conventional WS systolic array. As a result, DiP outperforms WS in terms of throughput, latency, area, and power consumption. Additionally, the analytical models for latency, throughput, FIFO overhead, and TFPU are developed for DiP and WS architectures. The analytical model shows that DiP architecture achieves up to 50% improvement in throughput and TFPU over WS counterparts. Moreover, hardware design space exploration is presented for both DiP and WS architectures using commercial 22nm technology, demonstrating power consumption savings of up to 19.95%, area savings of up to 8.12% at 1 GHz, and energy efficiency per area improvement of up to 2.02×. Furthermore, DiP is evaluated using various transformer workloads from widely used models such as GPT-2, GPT-3, BERT, BART, and LLaMA. DiP outperforms TPU-like architecture, achieving energy improvements ranging from 1.25× to 1.81× and latency improvements ranging from 1.03× to 1.49× across various transformer MHA and FFN workloads. DiP achieves a peak performance of 8.192 TOPS with energy efficiency of 9.548 TOPS/W. A comparison between state-of-the-art accelerators and DiP is presented, demonstrating DiP’s potential as a promising architecture for energy-efficient matrix multiplication acceleration. This paper is the foundation for DiP architecture and dataflow. Future work will focus on leveraging multiple DiP cores and deploying optimized transformers to enhance energy efficiency and performance.